The point A divides the join of P(-5,1) and Q(3,5) in the ratio K:1. Find the two values of k for which the area of triangle ABC, where B is (1,5) and C is (7,-2), is equal to 2 units.

Grade:11

1 Answers

AJIT AskiitiansExpert-IITD
13 years ago

Dear chanchal ,

let the co-ordinates of  A be (x ,y). then x =  3k-5/k+1  ,  y =5k+1/k+1 

now use the determinant formulae for calculating the area of triangle ,

we get ,

|14k-66/k+1| = 4

taking +4 and -4 separately ,

14k-66 =4k+4  gives k =7  and 14k-66 = -4k-4 gives , k= 31/9

Please feel free to post as many doubts on our discussion forum as you can. We are all IITians and here to help you in your IIT JEE preparation.


Now you can win exciting gifts by answering the questions on Discussion Forum. So help discuss any query on askiitians forum and become an Elite Expert League askiitian.

Win exciting gifts by answering the questions on Discussion Forum.